# Simultaneous Source Separation, Synchronization, Localization and Mapping for 6G Systems

### v2

@@ Abstract
mpslam is a promising approach for future 6G networks to jointly estimate the positions of transmitters and receivers together with the propagation environment. In cooperative multipath-based simultaneous localization and mapping (MP-SLAM), information collected by multiple mobile terminals is fused to enhance accuracy and robustness. Existing methods, however, typically assume perfectly synchronized base stations and orthogonal transmission sequences, rendering inter-BS interference at the MTs negligible. In this work, we relax these assumptions and address simultaneous source separation, synchronization, and mapping. A relevant example arises in modern 5G systems, where BSs employ muting patterns to mitigate interference, yet localization performance still degrades. We propose a novel BS-dependent data association and synchronization bias model, integrated into a joint Bayesian framework and inferred via the sum-product algorithm on a factor graph. The impact of joint synchronization and source separation is analyzed under various system configurations. Compared with state-of-the-art cooperative MP-SLAM assuming orthogonal and synchronized BSs, our statistical analysis shows no significant performance degradation. The proposed BS-dependent data association model constitutes a principled approach for classifying features by arbitrary properties, such as reflection order or feature type (scatterers versus walls).
@@ I Introduction
In contrast to conventional methods that treat multipath as interference, MP-SLAM exploits it as an additional source of information by modeling specular reflections as virtual anchors (VAs)—the mirror images of BSs at reflecting surfaces [3]. The joint estimation of MTs and virtual anchors thus enables accurate and robust localization even in obstructed line-of-sight conditions [5, 6, 7, 8, 9, 10].
@@ III Radio Signal Model and Channel Estimation
where ${s}(\tau)[\ell]=S^{(j)}(\ell\Delta){e}^{\big(i2\pi\ell\Delta(\tau+{b_{\text{bs}}}_{n}^{(j)}-{b_{\text{mt}}}_{n}^{(i)})\big)}$ with ${b_{\text{bs}}}_{n}^{(j)}~-~{b_{\text{mt}}}_{n}^{(i)}$ being the observed clock bias. ${N_{\text{bs}}}_{n}^{(j,i)}$} is the number of visible path (related to VAs), $S^{(j)}(f)$ represents the signal spectrum, which is sampled at $\ell=-(M-1)/2,\hskip 2.84526pt\ldots\hskip 2.84526pt,(M-1)/2$, $h$ represent the index for BS antennas [^2] In 5G systems, access to the impulse responses of individual antenna elements is typically unavailable. Instead, beamforming is performed at the BS, and spreading sequences of individual beams are embedded within dedicated resource blocks (RBs). The signal model in (III) can straightforwardly be reformulated in the beam space following [22], allowing for equivalent application of parametric channel estimation according to Section III-A. Note that we assume the spreading sequences corresponding to different beams to be encoded in RBs within the same PRS symbol. If this assumption is dropped, data association must be carried out individually for each beam, since the inter-BS interference pattern then differs across beams. , $h^{\prime}$ represents the index for MT antennas, and $c$ is the speed-of-light. The quantities $d_{n}^{(i,h^{\prime})}$ and $\psi_{n}^{(i,h^{\prime})}$ (respectively, $d^{(j,h^{\prime})}$ and $\psi^{(j,h^{\prime})}$) denote the propagation delay and angle associated with the individual antenna element located at position $\mathbf{p}_{n}^{(h,i)}$ (respectively, $\mathbf{p}^{(h,j)}$) relative to the center of gravity of the corresponding array. ${\tau}_{l,n}^{(j,i)}$, $\varphi^{(j,i)}_{l,n}$ and $\vartheta^{(j,i)}_{l,n}$ denote respectively delay, AOA and AOD of the MPCs (see Section II). The associated complex amplitude $\alpha_{l,n}^{(j,i)}\in\mathbb{C}$ is given by ${\alpha}_{l,n}^{(j,i)}={a}^{(j,i)}_{l,n}\hskip 0.85358pt\text{e}^{-j2\pi f_{\text{c}}{\tau}_{l,n}^{(j,i)}}c/(4\pi f_{\text{c}}{d}_{l,n}^{(j,i)})$ where ${a}^{(j,i)}_{l,n}\in\mathbb{C}$ is a reflection coefficient originating from all interactions of the radio signal with the radio equipment and the associated flat surfaces. The last term ${w}_{n}^{(j,i)}[\ell,h]$ aggregates the measurement noise. The receive signal samples are collected into the vector $\bm{r}^{(i)}_{n}=\big[\bm{r}_{n}^{(i)\hskip 0.85358pt\text{T}}[1,1]\hskip 2.84526pt\cdots\hskip 2.84526pt\bm{r}_{n}^{(i)\hskip 0.85358pt\text{T}}[H,1]\hskip 2.84526pt\bm{r}_{n}^{(i)\hskip 0.85358pt\text{T}}[1,2]\hskip 2.84526pt\cdots\hskip 2.84526pt$ $\bm{r}_{n}^{(i)\hskip 0.85358pt\text{T}}[H,H^{\prime}]\big]^{\text{T}}\in\mathbb{C}^{MHH^{\prime}}$.
@@ IV-A Measurement Model and New PVAs
where the means $d^{(k,i)}_{k,n}$, $\varphi^{(k,i)}_{k,n}$ and $\vartheta^{(k,i)}_{k,n}$ are defined in Sec. II, and the variances, derived from the Fisher information, are $\sigma^{(i)\,2}_{d_{k,n}}=c^{2}/(8\pi^{2}\beta_{\mathrm{bw}}^{2}{z_{\mathrm{u}}}^{(i)\,2}_{m,n})$, ${\sigma_{\varphi}}^{(k,i)\,2}_{k,n}=1/\!\big(8\pi^{2}{z_{\mathrm{u}}}^{(i)\,2}_{m,n}D^{2}(\varphi^{(k,i)}_{k,n})\big)$, and ${\sigma_{\vartheta}}^{(k,i)\,2}_{k,n}=1/\!\big(8\pi^{2}{z_{\mathrm{u}}}^{(i)\,2}_{m,n}D^{2}(\vartheta^{(k,i)}_{l,n})\big)$. Here, $\beta_{\mathrm{bw}}^{2}$ denotes the mean square bandwidth and $D^{2}(\cdot)$ the normalized squared array aperture [24, 21]. Similarly, for the BS-related LHF, we get
@@ IV-B Legacy PVAs and Sequential Update
New PVAs become legacy PVAs when the next measurements—either of the next MT or at the next time instance—are taken into account. In particular, the VA represented by the new VA state $\overline{\bm{\mathsfbr{y}}}^{(j,i^{\prime})}_{m^{\prime},n^{\prime}}$ introduced due to measurement $m^{\prime}$ of MT $i^{\prime}$ corresponding to BS $j$ at $n^{\prime}\hskip-0.85358pt\hskip-0.85358pt\leqslant\hskip-0.85358pt\hskip-0.85358ptn$ is represented by the legacy PVA state $\underline{\bm{\mathsfbr{y}}}^{(i)}_{k,n}=[\underline{\bm{\mathsfbr{p}}}^{(i)\hskip 0.85358pt\text{T}}_{\text{va},m}\hskip 2.84526pt\underline{\mathsfbr{r}}^{(i)}_{m}]^{\text{T}}$ at time $n$, i.e.,
@@ IV-C State Evolution
$f\big(\underline{\bm{p}}_{\text{va},k},\underline{r}_{k,n}\big|\bm{p}_{\text{va},k},r_{k,n-1}\hskip-0.85358pt\hskip-0.85358pt=\hskip-0.85358pt\hskip-0.85358pt0\big)=\hskip-0.85358pt\begin{cases}f_{\text{d}}\big(\underline{\bm{p}}_{\text{va},k}\big),&\!\!\!\underline{r}_{k,n}\!=\hskip-0.85358pt0\\[0.0pt] 0,&\!\!\!\underline{r}_{k,n}\!=\!1.\end{cases}$ (13)
To account for the sequential update, we define $f^{(i)}\big(\underline{\bm{y}}^{(i)}_{k,n}\big|{\bm{y}}^{(i-1)}_{k,n}\big)$ for $i\geqslant 2$ as
$f^{(i)}\big(\underline{\bm{p}}^{(i)}_{\text{va},k},\underline{r}^{(i)}_{k,n}\hskip 0.85358pt\big|{\hskip 0.85358pt{\bm{p}}^{(i-1)}_{\text{va},k},{r}^{(i-1)}_{k,n}=1}\big)$
$\hskip-85.35826pt=\hskip-0.85358pt\begin{cases}f_{\text{d}}\big(\underline{\bm{p}}^{(i)}_{\text{va},k}\big),&\!\!\!\underline{r}^{(i)}_{k,n}\!=\hskip-0.85358pt0\\[0.0pt] \delta\big(\underline{\bm{p}}^{(i)}_{\text{va},k}\hskip-0.85358pt-\hskip 0.85358pt{\bm{p}}^{(i-1)}_{\text{va},k}\big),&\!\!\!\underline{r}^{(i)}_{k,n}\!=\!1\end{cases}$ (14)
and
$f^{(i)}\big(\underline{\bm{p}}^{(i)}_{\text{va},k},\underline{r}^{(i)}_{k,n}\big|{{\bm{p}}^{(i-1)}_{\text{va},k},{r}^{(i-1)}_{k,n}\hskip-0.85358pt\hskip-0.85358pt=\hskip-0.85358pt\hskip-0.85358pt0}\big)$
$\hskip-56.9055pt=\hskip-0.85358pt\begin{cases}f_{\text{d}}\big(\underline{\bm{p}}^{(i)}_{\text{va},k}\big),&\!\!\!\underline{r}^{(i)}_{k,n}\!=\hskip-0.85358pt0\\[0.0pt] 0,&\!\!\!\underline{r}^{(i)}_{k,n}\!=\!1.\end{cases}$ (15)
@@ IV-D Data Association
Data assocation between measurements and PVAs is described in a redundant formulation by the PVA/PA-oriented association vector $\underline{\bm{\mathsfbr{a}}}_{n}^{(i)}\triangleq[\underline{\mathsfbr{a}}_{1,n}^{(i)}\hskip 0.85358pt\cdots\hskip 0.85358pt\underline{\mathsfbr{a}}_{J,n}^{(i)}\hskip 2.84526pt\underline{\mathsfbr{a}}_{J+1,n}^{(i)}\hskip 0.85358pt\cdots\hskip 0.85358pt\underline{\mathsfbr{a}}_{J+\mathsfbr{K}^{(i)}_{n},n}^{(i)}]^{\mathrm{T}}$ and by the measurement-oriented association vector $\overline{\bm{\mathsfbr{a}}}_{n}^{(i)}\triangleq[\overline{\mathsfbr{a}}_{1,n}^{(i)}\hskip 0.85358pt\cdots\hskip 0.85358pt\overline{\mathsfbr{a}}_{{\mathsfbr{M}}_{n}^{(i)},n}^{(i)}]^{\mathrm{T}}$. See [25, 23, 6, 17] for details.
@@ V-2 Confirmation of PVAs and State Estimation
We estimate the MT states $\bm{\mathsfbr{x}}^{(i)}_{n}$ (including $\mathsfbr{b}_{\text{mt}\,n}^{(i)}$), the BS biases $\mathsfbr{b}_{\text{bs}\,n}^{(j)}$, and the PVA states $\bm{\mathsfbr{p}}_{\text{va},k}$ from all measurements $\bm{z}_{1:n}$ up to time $n$ using the respective minimum mean-square error (MMSE) estimates [26, Ch. 4] $\hat{{\bm{x}}}^{(i)}_{n}=[\hat{\tilde{\bm{x}}}^{(i)}_{n}\hskip 2.84526pt\hat{\bm{v}}^{(i)}_{n}\hskip 2.84526pt\hat{b}_{\text{mt},n}^{(i)}]$ and $\hat{b}_{\text{bs}\,n}^{(j)}$, which are calculated by the expected values of the respective marginal posterior distributions $f(\bm{x}^{(i)}_{n}|\bm{z}_{1:n})$, $f({b}_{\text{bs},n}^{(j)}|\bm{z}_{1:n})$. Detection of the PVAs $k\!\in\!\{1,\dots,K_{n}\}$ relies on the posterior existence probabilities $p(r_{k,n}\!=\!1|\bm{z}_{1:n})=\int f(\bm{p}_{\text{va},k},r_{k,n}\!=\!1|\bm{z}_{1:n})\,\mathrm{d}\bm{p}_{\text{va},k}$ and the conditional posteriors $f(\bm{p}_{\text{va},k}|r_{k,n}\!=\!1,\bm{z}_{1:n})=f(\bm{p}_{\text{va},k},r_{k,n}\!=\!1|\bm{z}_{1:n})/p(r_{k,n}\!=\!1|\bm{z}_{1:n})$. A PVA is confirmed if $p(r_{k,n}\!=\!1|\bm{z}_{1:n})>p_{\mathrm{cf}}$ [27].
Since introducing new PVAs causes $K_{n}$ to grow indefinitely, a suboptimal pruning step is applied. A PVAs is pruned if $p(r_{k,n}\!=\!1|\bm{z}_{1:n})<p_{\text{pr}}$, except for k=1k\!=\!1 of each BS, which is always retained. Note that before pruining $\underline{j}(k)$ needs to be saved, since (9) becomes invalid. For existing PVAs, we determine the MMSE estimate $\hat{\bm{p}}_{\text{va},k}$ by calculating the expected value of $f(\bm{p}_{\text{va},k}|r_{k,n}\!=\!1,\bm{z}_{1:n})$. Since direct marginalization of (16) is intractable, sequential message passing via the SPA on a factor graph [28] provides efficient approximations (“beliefs”) of all marginal posteriors $f(\bm{x}^{(i)}_{n}|\bm{z}_{1:n})$, $f(b_{\text{bs},n}^{(j)}|\bm{z}_{1:n})$, $f(\bm{p}_{\text{va},k}|r_{k,n}\!=\!1,\bm{z}_{1:n})$, and $p(r_{k,n}|\bm{z}_{1:n})$, for all MT states $\bm{x}^{(i)}_{n}$, BS biases $b_{\text{bs},n}^{(j)}$, and PVAs $k$.

### v3

@@ Abstract
mpslam is a promising approach for future 6G networks to jointly estimate the positions of transmitters and receivers together with the propagation environment. In cooperative multipath-based simultaneous localization and mapping (MP-SLAM), information collected by multiple mobile terminals is fused to enhance accuracy and robustness. Existing methods, however, typically assume perfectly synchronized base stations and orthogonal transmission sequences, rendering inter-BS interference at the MTs negligible. In this work, we relax these assumptions and address simultaneous source separation, synchronization, and mapping. A relevant example arises in modern 5G systems, where BSs employ muting patterns to mitigate interference, yet localization performance still degrades. We propose a novel BS-dependent data association and synchronization bias model, integrated into a joint Bayesian framework and inferred via the sum-product algorithm on a factor graph. The impact of joint synchronization and source separation is analyzed under various system configurations. Compared with state-of-the-art cooperative MP-SLAM assuming orthogonal and synchronized BSs, our statistical analysis shows no significant performance degradation. The proposed BS-dependent data association model constitutes a principled approach for classifying features by arbitrary properties that persist over time, such as reflection order or feature type (scatter points versus walls).
@@ I Introduction
In contrast to conventional methods that treat multipath as interference, MP-SLAM exploits it as an additional source of information by modeling specular reflections as virtual anchors (VAs)—the mirror images of BSs at reflecting surfaces [3]. The joint estimation of MTs and virtual anchors thus enables accurate and robust localization even in obstructed line-of-sight conditions [5, 6, 7, 8, 9, 10, 11].
@@ III Radio Signal Model and Channel Estimation
where ${s}(\tau)[\ell]=S^{(j)}(\ell\Delta){e}^{\big(i2\pi\ell\Delta(\tau+{b_{\text{bs}}}_{n}^{(j)}-{b_{\text{mt}}}_{n}^{(i)})\big)}$ with ${b_{\text{bs}}}_{n}^{(j)}~-~{b_{\text{mt}}}_{n}^{(i)}$ being the observed clock bias. ${N_{\text{bs}}}_{n}^{(j,i)}$} is the number of visible path (related to VAs), $S^{(j)}(f)$ represents the signal spectrum, which is sampled at $\ell=-(M-1)/2,\hskip 2.84526pt\ldots\hskip 2.84526pt,(M-1)/2$, $h$ represent the index for BS antennas [^2] In 5G systems, access to the impulse responses of individual antenna elements is typically unavailable. Instead, beamforming is performed at the BS, and spreading sequences of individual beams are embedded within dedicated resource blocks (RBs). The signal model in (III) can straightforwardly be reformulated in the beam space following [23], allowing for equivalent application of parametric channel estimation according to Section III-A. Note that we assume the spreading sequences corresponding to different beams to be encoded in RBs within the same PRS symbol. If this assumption is dropped, data association must be carried out individually for each beam, since the inter-BS interference pattern then differs across beams. , $h^{\prime}$ represents the index for MT antennas, and $c$ is the speed-of-light. The quantities $d_{n}^{(i,h^{\prime})}$ and $\psi_{n}^{(i,h^{\prime})}$ (respectively, $d^{(j,h^{\prime})}$ and $\psi^{(j,h^{\prime})}$) denote the propagation delay and angle associated with the individual antenna element located at position $\mathbf{p}_{n}^{(h,i)}$ (respectively, $\mathbf{p}^{(h,j)}$) relative to the center of gravity of the corresponding array. ${\tau}_{l,n}^{(j,i)}$, $\varphi^{(j,i)}_{l,n}$ and $\vartheta^{(j,i)}_{l,n}$ denote respectively delay, AOA and AOD of the MPCs. The complex amplitude $\alpha_{l,n}^{(j,i)}\in\mathbb{C}$ is given by ${\alpha}_{l,n}^{(j,i)}={a}^{(j,i)}_{l,n}\hskip 0.85358pt\text{e}^{-j2\pi f_{\text{c}}{\tau}_{l,n}^{(j,i)}}c/(4\pi f_{\text{c}}{d}_{l,n}^{(j,i)})$ where ${a}^{(j,i)}_{l,n}\in\mathbb{C}$ is a reflection coefficient originating from all interactions of the radio signal with the radio equipment and the associated flat surfaces. The last term ${w}_{n}^{(j,i)}[\ell,h]$ aggregates the measurement noise. The according received signal vector is $\bm{r}^{(i)}_{n}=\big[\bm{r}_{n}^{(i)\hskip 0.85358pt\text{T}}[1,1]\hskip 2.84526pt\cdots\hskip 2.84526pt\bm{r}_{n}^{(i)\hskip 0.85358pt\text{T}}[H,1]\hskip 2.84526pt\bm{r}_{n}^{(i)\hskip 0.85358pt\text{T}}[1,2]\hskip 2.84526pt\cdots\hskip 2.84526pt$ $\bm{r}_{n}^{(i)\hskip 0.85358pt\text{T}}[H,H^{\prime}]\big]^{\text{T}}\in\mathbb{C}^{MHH^{\prime}}$.
@@ IV-A Measurement Model and New PVAs
where the means $d^{(k,i)}_{k,n}$, $\varphi^{(k,i)}_{k,n}$ and $\vartheta^{(k,i)}_{k,n}$ are defined in Section II, and the variances, derived from the Fisher information, are $\sigma^{(i)\,2}_{d_{k,n}}=c^{2}/(8\pi^{2}\beta_{\mathrm{bw}}^{2}{z_{\mathrm{u}}}^{(i)\,2}_{m,n})$, ${\sigma_{\varphi}}^{(k,i)\,2}_{k,n}=1/\!\big(8\pi^{2}{z_{\mathrm{u}}}^{(i)\,2}_{m,n}D^{2}(\varphi^{(k,i)}_{k,n})\big)$, and ${\sigma_{\vartheta}}^{(k,i)\,2}_{k,n}=1/\!\big(8\pi^{2}{z_{\mathrm{u}}}^{(i)\,2}_{m,n}D^{2}(\vartheta^{(k,i)}_{l,n})\big)$. Here, $\beta_{\mathrm{bw}}^{2}$ denotes the mean square bandwidth and $D^{2}(\cdot)$ the normalized squared array aperture [25, 22]. Similarly, for the BS-related LHF, we get
@@ IV-B Legacy PVAs and Sequential Update
New PVAs become legacy PVAs when the next measurements, either of the next MT or at the next time instance, are taken into account. In particular, the VA represented by the new VA state $\overline{\bm{\mathsfbr{y}}}^{(j,i^{\prime})}_{m^{\prime},n^{\prime}}$ introduced due to measurement $m^{\prime}$ of MT $i^{\prime}$ corresponding to BS $j$ at $n^{\prime}\hskip-0.85358pt\hskip-0.85358pt\leqslant\hskip-0.85358pt\hskip-0.85358ptn$ is represented by the legacy PVA state $\underline{\bm{\mathsfbr{y}}}^{(i)}_{k,n}=[\underline{\bm{\mathsfbr{p}}}^{(i)\hskip 0.85358pt\text{T}}_{\text{va},m}\hskip 2.84526pt\underline{\mathsfbr{r}}^{(i)}_{m}]^{\text{T}}$ at time $n$, i.e.,
@@ IV-C State Evolution
$f\big(\underline{\bm{p}}_{\text{va},k},\underline{r}_{k,n}\big|\bm{p}_{\text{va},k},0\big)=\hskip-0.85358pt\begin{cases}f_{\text{d}}\big(\underline{\bm{p}}_{\text{va},k}\big),&\!\!\!\underline{r}_{k,n}\!=\hskip-0.85358pt0\\[0.0pt] 0,&\!\!\!\underline{r}_{k,n}\!=\!1.\end{cases}$ (13)
To account for the sequential update, we define $f^{(i)}\big(\underline{\bm{y}}^{(i)}_{k,n}\big|{\bm{y}}^{(i-1)}_{k,n}\big)$ for $i\geqslant 2$ and $\underline{r}^{(i)}_{k-1,n}=1$ as
$\hskip-11.38109ptf^{(i)}\big(\underline{\bm{p}}^{(i)}_{\text{va},k},\underline{r}^{(i)}_{k,n}\hskip 0.85358pt\big|{\hskip 0.85358pt{\bm{p}}^{(i-1)}_{\text{va},k},1}\big)\hskip-0.85358pt=\hskip-0.85358pt\hskip-0.85358pt\begin{cases}f_{\text{d}}\big(\underline{\bm{p}}^{(i)}_{\text{va},k}\big),&\!\!\!\underline{r}^{(i)}_{k,n}\!=\hskip-0.85358pt0\\[0.0pt] \delta\big(\underline{\bm{p}}^{(i)}_{\text{va},k}\hskip-0.85358pt-\hskip 0.85358pt{\bm{p}}^{(i-1)}_{\text{va},k}\big),&\!\!\!\underline{r}^{(i)}_{k,n}\!=\!1\end{cases}$ (14)
and for $\underline{r}^{(i)}_{k-1,n}=0$
$f^{(i)}\big(\underline{\bm{p}}^{(i)}_{\text{va},k},\underline{r}^{(i)}_{k,n}\big|{{\bm{p}}^{(i-1)}_{\text{va},k},0}\big)=\hskip-0.85358pt\begin{cases}f_{\text{d}}\big(\underline{\bm{p}}^{(i)}_{\text{va},k}\big),&\!\!\!\underline{r}^{(i)}_{k,n}\!=\!0\\[0.0pt] 0,&\!\!\!\underline{r}^{(i)}_{k,n}\!=\!1.\end{cases}$ (15)
@@ IV-D Data Association
Data association between measurements and PVAs is described in a redundant formulation by the PVA/PA-oriented association vector $\underline{\bm{\mathsfbr{a}}}_{n}^{(i)}\triangleq[\underline{\mathsfbr{a}}_{1,n}^{(i)}\hskip 0.85358pt\cdots\hskip 0.85358pt\underline{\mathsfbr{a}}_{J,n}^{(i)}\hskip 2.84526pt\underline{\mathsfbr{a}}_{J+1,n}^{(i)}\hskip 0.85358pt\cdots\hskip 0.85358pt\underline{\mathsfbr{a}}_{J+\mathsfbr{K}^{(i)}_{n},n}^{(i)}]^{\mathrm{T}}$ and by the measurement-oriented association vector $\overline{\bm{\mathsfbr{a}}}_{n}^{(i)}\triangleq[\overline{\mathsfbr{a}}_{1,n}^{(i)}\hskip 0.85358pt\cdots\hskip 0.85358pt\overline{\mathsfbr{a}}_{{\mathsfbr{M}}_{n}^{(i)},n}^{(i)}]^{\mathrm{T}}$. See [26, 24, 6, 18] for details.
@@ V-2 Confirmation of PVAs and State Estimation
We estimate the MT states $\bm{\mathsfbr{x}}^{(i)}_{n}$ (including $\mathsfbr{b}_{\text{mt}\,n}^{(i)}$), the BS biases $\mathsfbr{b}_{\text{bs}\,n}^{(j)}$, and the PVA states $\bm{\mathsfbr{p}}_{\text{va},k}$ from all measurements $\bm{z}_{1:n}$ up to time $n$ using the respective minimum mean-square error (MMSE) estimates [27, Ch. 4] $\hat{{\bm{x}}}^{(i)}_{n}=[\hat{\tilde{\bm{x}}}^{(i)}_{n}\hskip 2.84526pt\hat{\bm{v}}^{(i)}_{n}\hskip 2.84526pt\hat{b}_{\text{mt},n}^{(i)}]$ and $\hat{b}_{\text{bs}\,n}^{(j)}$, which are calculated by the expected values of the respective marginal posterior distributions $f(\bm{x}^{(i)}_{n}|\bm{z}_{1:n})$, $f({b}_{\text{bs},n}^{(j)}|\bm{z}_{1:n})$. Detection of the PVAs $k\!\in\!\{1,\dots,K_{n}\}$ relies on the posterior existence probabilities $p(r_{k,n}\!=\!1|\bm{z}_{1:n})=\int f(\bm{p}_{\text{va},k},r_{k,n}\!=\!1|\bm{z}_{1:n})\,\mathrm{d}\bm{p}_{\text{va},k}$ and the conditional posteriors $f(\bm{p}_{\text{va},k}|r_{k,n}\!=\!1,\bm{z}_{1:n})=f(\bm{p}_{\text{va},k},r_{k,n}\!=\!1|\bm{z}_{1:n})/p(r_{k,n}\!=\!1|\bm{z}_{1:n})$. A PVA is confirmed if $p(r_{k,n}\!=\!1|\bm{z}_{1:n})>p_{\mathrm{cf}}$ [28].
Since introducing new PVAs causes $K_{n}$ to grow indefinitely, a suboptimal pruning step is applied. A PVAs is pruned if $p(r_{k,n}\!=\!1|\bm{z}_{1:n})<p_{\text{pr}}$, except for k=1k\!=\!1 of each BS, which is always retained. Note that before pruining $\underline{j}(k)$ needs to be saved, since (9) becomes invalid. For existing PVAs, we determine the MMSE estimate $\hat{\bm{p}}_{\text{va},k}$ by calculating the expected value of $f(\bm{p}_{\text{va},k}|r_{k,n}\!=\!1,\bm{z}_{1:n})$. Since direct marginalization of (16) is intractable, sequential message passing via the SPA on a factor graph [29] provides efficient approximations (“beliefs”) of all marginal posteriors $f(\bm{x}^{(i)}_{n}|\bm{z}_{1:n})$, $f(b_{\text{bs},n}^{(j)}|\bm{z}_{1:n})$, $f(\bm{p}_{\text{va},k}|r_{k,n}\!=\!1,\bm{z}_{1:n})$, and $p(r_{k,n}|\bm{z}_{1:n})$, for all MT states $\bm{x}^{(i)}_{n}$, BS biases $b_{\text{bs},n}^{(j)}$, and PVAs $k$. The runtime of the proposed algorithm scales linearly in the number of PVAs, measurements, MTs and particles, i.e, $\mathcal{O}(JK_{n-1}^{(i)}M_{n}^{(i)}P)$. The number of PVAs, which increases over time $n$ and MT index $i$ according to (11), can become large. Compared to state-of-the-art cooperative MP-SLAM [18] the proposed algorithm involves an additional factor $J$, since a new PVA is generated for each anchor. This is counteracted by pruning.